please answer my question: Andy's weight is 5 kilograms less than twice his brother's. Together they both weigh 100 kilograms. What are their weights?
Found 2 solutions by checkley77, Fombitz:
Answer by checkley77(12844)

Let x be the brother's weight.
Then 2x-5 is Andy's weight.
x+(2x-5)=100
3x-5=100
3x=105
x=105/3
x=35 pounds is the brother's weight.
2*35-5=70-5=65 pounds is Andy's weight.
Proof:
35+65=100
100=100

Answer by Fombitz(32388)